OpenStudy (anonymous):

find the axis of symmetry of the parabola given by y=x^2 - 14x + 3

OpenStudy (anonymous):

Axis of Symmetry is the x value of the vertex

OpenStudy (anonymous):

\[Vertex = (\frac{-b}{2a}, f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

2x - 14 = 0 2x = 14 axis of symmetry x = 7